Can you please complete the questions about the supplier will only start to supply t-shirts when a price greater then k50 per unit is available.
1
Expert's answer
2021-09-23T12:54:27-0400

complete question

A supplier will only start to supply T shirts when a prize when a price greater than ₹500 per unit is available. He or she will then increase the output by 2 units for every ₹50 per unit increase in price. Write the equation of the supply function.

solution

Let the supplier will start to supply T shirts when a price is ₹500 per unit. He or she will then increase the output by 2 units for every ₹50 per unit increase in price. Then, we can write the supply function as follows:



"P_s=mQ+b,"

here, "m" is the slope of the supply curve.

We can find the slope as follows:



"m=\dfrac{\Delta P}{\Delta Q}=\dfrac{50}{2}=25."

To find coefficient "b" let's plug into the supply function the quantity supplied 2 and corresponding price ₹550 (₹500+₹50):



"550=25\cdot2+b,""b=500."

Finally, we can write the equation of the supply function:



"P_s=25Q+500."